OpenStudy (anonymous):

solve -10over13 = -2 1over2 b will give medal to best answer.

OpenStudy (anonymous):

is that 21?

OpenStudy (anonymous):

no, i think it is a fraction .

OpenStudy (anonymous):

well i got 20over273 .-.

OpenStudy (anonymous):

lol, i don't think that is right. it is a choice question. 1. b = -4/13 2. b=1 12/13 3. b= 4/13 4. b = -1 12/13

OpenStudy (anonymous):

@*amy*

OpenStudy (anonymous):

Agghhh I cant get it but its most likely not 1 or 3 because the negatives